### v1

# Image Complexity-Aware Adaptive Retrieval for Efficient Vision-Language Models

Abstract
Vision transformers in vision-language models apply uniform computational effort across all images, expending 175.33 GFLOPs (ViT-L/14) whether analysing a straightforward product photograph or a complex street scene. We propose ICAR (Image Complexity-Aware Retrieval), which enables vision transformers to use less compute for simple images whilst processing complex images through their full network depth. The key challenge is maintaining cross-modal alignment: embeddings from different processing depths must remain compatible for text matching. ICAR solves this through dual-path training that produces compatible embeddings from both reduced-compute and full-compute processing. This maintains compatibility between image representations and text embeddings in the same semantic space, whether an image exits early or processes fully. Unlike existing two-stage approaches that require expensive reranking, ICAR enables direct image-text matching without additional overhead. To determine how much compute to use, we develop ConvNeXt-IC, which treats image complexity assessment as a classification task. By applying modern classifier backbones rather than specialised architectures, ConvNeXt-IC achieves state-of-the-art performance with 0.959 correlation with human judgement (Pearson) and 4.4× speedup. Evaluated on standard benchmarks augmented with real-world web data, ICAR achieves 20% practical speedup while maintaining category-level performance and 95% of instance-level performance, enabling sustainable scaling of vision-language systems.
1 Introduction
Modern search engines use Vision Transformers (ViT) within vision-language models to match images with text queries, enabling applications from web image search to e-commerce product matching. ViTs process images through multiple layers in sequence, with each layer adding more detailed understanding of the image [raghu2021vision].
Standard ViT systems have a fundamental inefficiency: they use the same amount of compute for all images, regardless of the complexity of the image. For example, a 24-layer ViT-L/14 uses the same 175.33 GFLOPs [ilharco_gabriel_2021_5143773] of compute whether it’s processing a simple product photo or a complex street scene. At web scale, where major platforms process billions of images daily [GoogleIO2024], this uniform processing can lead to a substantial waste of computational resources. The logical solution is adaptive computation: allowing simple images to use less compute while complex images use more compute.
Adaptive computation can be achieved through early exit: given that ViTs process images through a sequence of layers that progressively refine extracted features, simple images can stop processing at intermediate layers [raghu2021vision], while complex images continue through the full network. Adaptive computation for vision-language models faces a fundamental technical challenge: maintaining cross-modal alignment when embeddings are produced from different processing depths. While adaptive computation has been extensively studied for single-modality tasks [teerapittayanon2016branchynet], the adaptive vision-language field remains underexplored. Specifically, no existing work addresses adaptive computation for image-text retrieval tasks.
We address this by proposing ICAR (Image Complexity-Aware Retrieval), which leverages image complexity to determine the processing path for each image (Figure 1). ICAR introduces dual-path training that produces compatible embeddings from both early reduced compute and full processing paths. This enables simple images to use fewer layers whilst complex images process through the full network, all while maintaining cross-modal alignment through unified embedding spaces, eliminating the reranking overhead that existing approaches have. Evaluated on standard benchmarks (Flickr30K [young2014flickr30k]/MS-COCO [lin2014microsoft]) augmented with 100K real-world samples from LAION-5B, ICAR achieves 20% practical speedup while maintaining category-level performance and 95% of instance-level performance.
To enable ICAR’s routing decisions, we develop ConvNeXt-IC (ConvNeXt-Image Complexity). While existing methods treat complexity assessment as a representation learning problem requiring specialised architectures, we instead recognise it as a classification task. Therefore, by using modern classification backbones pretrained on ImageNet [deng2009imagenet] rather than self-supervised representation learning approaches, ConvNeXt-IC achieves state-of-the-art performance (0.959 Pearson correlation, 4.4 speedup). We validate across both established academic benchmarks (IC9600) and real-world web data (LAION), demonstrating that our approach generalises beyond controlled settings. The following are the contributions of the paper:
First adaptive computation method for image-text retrieval: We introduce ICAR, which addresses the cross-modal alignment challenge through dual-path training that produces compatible embeddings from different processing depths, enabling direct text matching without reranking overhead.
Reconceptualising complexity assessment as classification: We demonstrate that image complexity assessment is better suited as a classification task, rather than representation learning. By treating it as classification with modern backbones (ConvNeXt-V2) instead of building specialised complexity assessment architectures, we achieve state-of-the-art performance (0.959 Pearson correlation) with 4.4 speedup over existing representation learning methods.
Differential computational requirements: We observe that category-level and instance-level retrieval have different computational needs: category matching requires only coarse semantic features while instance matching demands fine-grained discrimination, enabling targeted efficiency optimisations with 20% practical speedup.
This paper is organised as follows: Section 2 discusses related work in image complexity assessment and adaptive computation. Section 3 presents our ConvNeXt-IC complexity detection method with validation on established benchmarks and real-world data. Section 4 details the ICAR architecture and dual-path training approach. Section 5 presents experimental results and analysis of retrieval performance and efficiency gains. Finally, Section 6 concludes with key findings and future directions.
2 Related Work
2.1 Image Complexity Assessment
Image complexity assessment determines how difficult an image is for humans to understand or describe [snodgrass1980standardized]. Psychology research identified key factors: object count, spatial relationships, visual diversity, and background richness [olivia2004identifying, forsythe2009visual, purchase2012predicting]. Early computational approaches used hand-crafted features like entropy and compression ratios [stamps2002entropy, machado2015computerized, rosenholtz2007measuring], but these worked only on small datasets.
Current deep learning methods treat complexity assessment as a representation learning problem. ICNet [feng2022ic9600] uses a dual-branch architecture combining detail and context pathways with spatial attention modules. ICCORN [guo2023image] applies ordinal regression to ICNet features, explicitly modelling the ordinal structure of complexity ratings. MICM [li2025unlocking] introduces motion-inspired assessment using image-to-video generation with hierarchical alignment losses. CLICv2 [liu2025clicv2] uses self-supervised pretraining on complexity-related visual patterns.
Related image quality assessment methods explore similar perceptual challenges but focus on quality rather than complexity. HyperIQA [su2020blindly] employs content-adaptive hyper networks for quality assessment, while CLIPIQA [wang2022exploring] leverages vision-language representations from CLIP [radford2021learning]. TOPIQ [chen2024topiq] adopts a top-down framework modelling human visual perception characteristics.
These methods follow a common pattern: start with classification baselines, then add complexity-specific architectures for incremental improvements. However, they test only against outdated networks like AlexNet [krizhevsky2012imagenet] (2012) and ResNet [he2016deep] (2016), leaving a large gap unexplored. This raises an open question whether architectural specialisation provides genuine benefits, or whether stronger classification backbones would close the performance gap without the need for additional components. We address this gap using ConvNeXt-V2 (2023) to test whether modern classification architectures outperform specialised complexity assessment designs.
2.2 Adaptive Computation in Machine Learning
Without adaptive computation, a vision-language model such as OpenCLIP’s ViT-L/14 uses 175.33 GFLOPs for each sample, regardless of whether the sample is challenging or easy to represent [ilharco_gabriel_2021_5143773]. The vision encoder consumes 162.03 GFLOPs while text processing requires only 13.3 GFLOPs, yet both apply the same computational effort to all inputs. This uniform processing creates inefficiencies where simple product photos receive the same 24-layer transformer processing as complex street scenes with dozens of objects [dosovitskiy2020image].
Single-modality methods have explored adaptive computation to address this inefficiency by adjusting processing based on input difficulty. BranchyNet [teerapittayanon2016branchynet] introduced early exits with classifiers at intermediate layers, enabling confident predictions to exit early while difficult cases continue through all layers. Several methods adapt early exits for Vision Transformers through patch processing, token removal, and routing strategies [yin2022avit, chen2023cf, rao2021dynamicvit, bolya2023token, riquelme2021scaling].
These vision-only methods cannot handle vision-language models because they ignore cross-modal alignment requirements. Image-text retrieval needs embeddings from different processing depths to remain compatible in the same semantic space for text matching: a fundamental challenge that existing adaptive methods cannot address.
Limited work exists on adaptive computation for vision-language models. DeeCap [fei2022deecap] and MuE [tang2023you] target generative tasks, while HiVLP [chen2022hivlp] uses hierarchical features for retrieval but requires re-ranking due to separate embedding spaces. No existing work enables direct single-stage retrieval with embedding compatibility across processing depths.
At web scale, where platforms process billions of images daily, efficiency gains have immediate infrastructure impact through reduced energy costs and faster response times. No existing work addresses adaptive computation for image-text retrieval, leaving this critical efficiency challenge unsolved.
3 Image Complexity Detection Method
This section presents and validates our proposed ConvNeXt-IC image complexity detection method. This method acts as a component for our proposed image-text retrieval method ICAR, the integration of which is discussed in Section 4.
3.1 Proposed Method: ConvNeXt-IC
To adapt image-text retrieval models to use more or less compute based on image complexity, we need automatic complexity assessment. Real-world web data exhibits a wide range of complexity levels. Figure 3 illustrates this with representative cases from the LAION-5B dataset [schuhmann2022laion].
These examples illustrate the range of complexity that methods must handle. Existing approaches treat complexity assessment as a representation learning problem, building specialised architectures to capture complexity patterns [feng2022ic9600, guo2023image, li2025unlocking]. However, the core challenge is discriminating between different levels of visual complexity. We therefore treat complexity assessment as a classification problem, using efficient classification backbones instead of specialised complexity architectures.
We therefore fine-tune ConvNeXt-V2-N (224x224) [woo2023convnext], the smallest of the ConvNeXt-V2 ImageNet-22K [deng2009imagenet] family of foundation classification models. This choice prioritises efficiency for real-time routing decisions while maintaining the discriminative power needed for accurate complexity assessment. We implement ConvNeXt-IC with two configurations: (1) As a regression model (using a single linear layer predicting continuous scores); and (2) as a binary classification model (using a lightweight two-layer MLP with cross-entropy loss). Both configurations are fine-tuned end-to-end using standard data augmentation (random crops, horizontal flips, colour jittering) on the benchmark datasets.
3.2 Validation on IC9600 Benchmark
To validate our approach, we evaluate ConvNeXt-IC on IC9600 [feng2022ic9600] in its regression head setting, the established benchmark for image complexity assessment. IC9600 contains 9,600 images labelled with complexity scores from 1 (simple) to 5 (complex). We measure performance using standard regression metrics: Pearson Correlation Coefficient (PCC) and Spearman Rank Correlation Coefficient (SRCC) assess how well predicted scores correlate with human ratings, while Root Mean Square Error (RMSE) measures prediction accuracy. We also report inference speed (images/second) to evaluate computational efficiency. Table 1 shows our results compared to existing methods.
| Method | PCC | SRCC | RMSE | img/s |
| HyperIQA | 0.935 | 0.926 | 0.204 | 61 |
| CLIPIQA | 0.898 | 0.897 | 0.078 | 83 |
| TOPIQ | 0.944 | 0.938 | 0.049 | 125 |
| CLICv2 | 0.933 | 0.927 | - | - |
| ICNet | 0.949 | 0.945 | 0.053 | 397 |
| MICM | 0.953 | 0.943 | 0.049 | 0.01 |
| ICCORN | 0.955 | 0.951 | - | - |
| ConvNeXt-IC (Ours) | 0.959 | 0.956 | 0.045 | 1744 |
ConvNeXt-IC achieves state-of-the-art performance: 0.959 PCC and 0.956 SRCC, outperforming the previous best method ICCORN (0.955 PCC, 0.951 SRCC) while delivering 1,744 images/second throughput. Most notably, ConvNeXt-IC surpasses methods that were specifically architecturally designed for image complexity detection, like ICNet (0.949 PCC) and MICM (0.953 PCC). This demonstrates that strong, correctly chosen backbones which are then fine-tuned can outperform specialised architectures, providing another instance of the current trend in machine learning where data and compute together can outperform sophisticated architectural design [halevy2009unreasonable, zhai2022scaling].
3.3 Domain Shift Spot Check on LAION Subset
While IC9600 provides a controlled validation benchmark, its images are sourced from curated academic datasets. To verify that ConvNeXt-IC generalises to the noisy, real-world web content that ICAR is targeting, we conduct a domain shift spot check on LAION data.
We sample 3,000 images from LAION-5B [schuhmann2022laion] and manually annotate them into complexity classes of ‘complex’ and ‘simple’. The annotation process naturally resulted in the classes being distributed as 49.5% simple and 50.5% complex. Note that, this annotation process was conducted solely by the lead researcher of this paper. Therefore, while the primary validation of ConvNeXt-IC was conducted on IC9600 as discussed in the previous subsection, experiments on this data are instead best interpreted as a domain shift spot check. These annotations are made publicly available in the project’s repository at https://github.com/MikelWL/ICAR.
The 3,000 annotated images are split into 2,100/450/450 (train/val/test) sets. For comparison, we use ICNet [feng2022ic9600] as the baseline; the most established complexity assessment method with viable inference speed for practical deployment among existing approaches. Given this is a binary ‘complex’ or ‘simple’ classification task, we evaluate ConvNeXt-IC in its classification head setting. Table 2 shows the comparative performance.
| Method | Precision | Recall | F1 | ROC AUC | img/s |
| ICNet | 0.757 | 0.741 | 0.749 | 0.845 | 397 |
| ConvNeXt-IC | 0.869 | 0.862 | 0.865 | 0.944 | 1744 |
ConvNeXt-IC achieves 0.865 F1-score and 0.944 ROC AUC, outperforming ICNet (0.749 F1, 0.845 ROC AUC) with its 4.4 efficiency advantage (1,744 vs 397 images/second). These results support ConvNeXt-IC’s effective generalisation from curated benchmarks to real-world web distributions, supporting our approach for ICAR’s deployment requirements.
4 Proposed ICAR Architecture
4.1 Model Design and Dual-Path Training
ICAR integrates ConvNeXt-IC complexity routing with adaptive ViT-L/14 processing. Figure 4 shows the architecture: simple images exit early at layers 8, 12, 16, or 20, while complex images process through all 24 layers. The approach applies dual-path training to produce compatible embeddings from both paths in the same semantic space, addressing the cross-modal alignment challenge in vision-language retrieval.
Architecture: ICAR consists of (1) ConvNeXt-IC for routing decisions (2.45 GFLOPs overhead), (2) ViT-L/14 with early exit points, and (3) identical projection heads ensuring embedding compatibility. Early exit at layer 8 saves 44% computation (108 GFLOPs), while later exits provide efficiency-quality trade-offs. Text processing remains unchanged at 13.3 GFLOPs.
Dual-Path Training: During training, each image-text pair passes through both the complexity-determined early exit path and the full processing path, creating two vision embeddings that must both align with the corresponding text embedding. We optimise a combined loss:
where and are CLIP contrastive losses from respective paths, and weights both paths equally. This creates multi-scale supervision that prevents early layers from over-specialising while ensuring both processing paths learn compatible representations within the same semantic space.
This unified approach eliminates the storage overhead and computational complexity of reranking methods, which require separate embedding storage and two-stage retrieval pipelines. All embeddings, regardless of processing depth, remain directly comparable using cosine similarity without additional processing steps.
4.2 Experimental Setup
We evaluate four ICAR variants (ICAR-8/24, ICAR-12/24, ICAR-16/24, ICAR-20/24) indicating early exit and total layers, using ICC complexity-based routing where simple images use early exits while complex images process through all layers. We train on MS-COCO [lin2014microsoft] (82K) and Flickr30K [young2014flickr30k] (29K), then evaluate on both datasets augmented with 100K LAION samples to test web-scale performance. As a web-scraped dataset, LAION provides the rich complexity distribution needed to evaluate adaptive computation under realistic deployment conditions.
ICAR models fine-tune from OpenCLIP ViT-L/14 using AdamW [loshchilov2019decoupled] with differential learning rates (1e-6 backbone, 1e-4 projection heads) for 10 epochs. We measure both retrieval quality and computational efficiency: instance-level retrieval uses Recall@k (exact image-caption matching), category-level retrieval uses mAP@k (broader semantic matching), and efficiency uses practical throughput on A100 GPUs.
5 Results and Analysis
5.1 Retrieval Performance and Key Insights
| Instance Retrieval - Text to Image - R@k | |||||||
| Flickr + LAION | COCO + LAION | ||||||
| @1 | @5 | @10 | @1 | @5 | @10 | GFLOPs | |
| ICAR-20/24 | 71.1 | 89.8 | 93.7 | 47.4 | 72.7 | 81.1 | 164.41 |
| ICAR-16/24 | 69.5 | 88.9 | 93.6 | 47.1 | 72.2 | 81.2 | 151.05 |
| ICAR-12/24 | 68.5 | 88.8 | 92.9 | 45.1 | 70.2 | 79.3 | 137.68 |
| ICAR-8/24 | 67.4 | 87.6 | 92.2 | 42.9 | 68.3 | 77.4 | 124.31 |
| ViT-L/14 (FT) | 71.0 | 89.9 | 93.8 | 48.5 | 73.5 | 81.8 | 175.33 |
| Category Retrieval - Text to Image - mAP@k | |||||||
| Flickr + LAION | COCO + LAION | ||||||
| @10 | @50 | @100 | @10 | @50 | @100 | GFLOPs | |
| ICAR-20/24 | 83.8 | 71.7 | 64.2 | 92.9 | 87.6 | 84.7 | 164.41 |
| ICAR-16/24 | 83.4 | 71.4 | 64.0 | 93.1 | 87.8 | 84.8 | 151.05 |
| ICAR-12/24 | 83.4 | 70.5 | 63.4 | 92.8 | 87.4 | 84.4 | 137.68 |
| ICAR-8/24 | 82.2 | 69.7 | 62.7 | 93.1 | 87.9 | 84.9 | 124.31 |
| ViT-L/14 (FT) | 84.3 | 70.2 | 64.1 | 92.6 | 86.2 | 82.2 | 175.33 |
Instance vs category: Instance and category retrieval have different computational requirements. Instance-level retrieval shows expected performance decline with early exits; for ICAR-8/24, the aggregate recall scores across Flickr+LAION and COCO+LAION is 95.0% of the baseline111Performance retention calculated using aggregate RSUM [li2022image] methodology: sum of all ICAR-8/24 retrieval scores divided by sum of baseline scores across both datasets.. For category-level retrieval, ICAR variants generally maintain or exceed baseline performance with small, dataset-specific differences: on COCO+LAION, mAP@10 can be marginally higher (e.g., 93.1 vs 92.6 baseline), whereas on Flickr+LAION, ICAR variants remain within 1–2 points of baseline across mAP@k.
Hypothesis validation: The results support our hypothesis about the progressive nature of transformer processing: early layers capture coarse semantic features adequate for category matching, while deeper layers refine instance-level discriminative details. Our dual-path training acts as implicit regularisation, creating more robust categorical representations through multi-scale supervision.
Regularisation effect: The maintained category-level performance is consistent with dual-path training acting as an implicit regularisation technique. Multi-scale supervision from both early and full path losses helps learn robust categorical representations. Early layers focus on coarse semantic features sufficient for category matching, while deeper layers refine instance-level details.
5.2 Efficiency Analysis and Impact
ICAR achieves 20% practical speedup (ICAR-8/24: 1.195, 120.0 vs 100.4 images/second) with GFLOP-estimated potential up to 1.41. The gap between estimated and measured performance requires production-grade optimisations to fully realise GFLOP reduction potential. Figure 6(a) compares estimated versus measured A100 speedups, while Figure 6(b) shows differential impact: category-level retrieval maintains or exceeds baseline (ICAR-8/24: 100.1% retention), whereas instance-level shows expected degradation. This supports ICAR’s strategy: category matching requires only coarse semantic features from intermediate layers, while instance matching demands fine-grained features from deeper processing.
Production-Scale Impact: If efficiency improvements such as ICAR’s 20% were implemented at big tech production scale, processing Google Photos’ 6 billion daily images [GoogleIO2024] could save 667 A100 GPU-hours daily and 133,640 kWh annually, equivalent to eliminating 16.6 metric tonnes of CO2 emissions222Projected calculation: Google Photos processes 6 billion images daily [GoogleIO2024]. Production-grade optimisations achieve 3-7× speedups [NVIDIADocsOptimization], suggesting 500 images/second on A100 GPUs. Daily processing requires 3,333 GPU-hours (6 billion ÷ 500 ÷ 3,600 seconds/hour). With A100’s 500W consumption and datacenter PUE of 1.10 [GoogleDataCenters], annual consumption is 668,002 kWh. ICAR’s 20% reduction could save 667 GPU-hours daily and 133,640 kWh annually. CO2 calculation uses UK grid intensity [CarbonBrief2024]..
6 Conclusion
This paper presented ICAR (Image Complexity-Aware Retrieval), the first adaptive computation method for image-text retrieval that addresses the cross-modal alignment challenge through unified embedding spaces. At the core of ICAR is ConvNeXt-IC, a lightweight image complexity classifier that serves as the routing mechanism: it predicts whether an image is simple or complex and routes simple images to early exits while complex images use full processing. During training, both the early-exit and full paths are jointly supervised via dual-path contrastive learning so that all vision embeddings remain compatible with the text embedding in a single unified space. By recognising complexity assessment as a classification task, ConvNeXt-IC achieves state-of-the-art performance (0.959 Pearson correlation, 4.4 speedup), and ICAR’s dual-path training enables 20% practical speedup with maintained retrieval quality.
Future work could enhance the approach through dynamic exit selection based on intermediate confidence scores, end-to-end integration of complexity assessment within the vision-language learning objective, and production-grade optimisations to bridge the gap between GFLOP-estimated (1.41×) and measured (1.195×) speedup results.
6.0.1 Acknowledgements
To facilitate reproducibility and future research, all code, trained model weights, and the annotated LAION complexity dataset are publicly available at https://github.com/MikelWL/ICAR.
[…]
The authors have no competing interests to declare that are relevant to the content of this article.